40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12299     Entry time: Wed Jul 13 15:35:56 2016
Author: Lydia 
Type: Update 
Category: General 
Subject: Vent Progress - ETMY repositioned and removed 

[Lydia, Johannes]

Took photos to document the original OSEM orientation and wrote down the serial numbers for each position. We removed the OSEMs, moved the suspension to the accessible side of the table and took out the optic, which was brought to the clean room to have the magnets reglued. The ETMY chamber is now closed up with the OSEMs and clamps inside on the table, and should not need to be reopened until the magnets have been reattached. 

ELOG V3.1.3-